a gardener has 79 feet of fencing to be used to enclose a rectangular garden that has a border 2 feet wide surrounding it. If the length and width of the garden were to be the same, what would be the dimensions of the garden?
Answer by mananth(16946)

Since the length and width are equal it is a square.
let the side be x
there is 2 feet border on each side
so the side including border will x+4 _________ 2 on each side
4(x+4)= 79
4x+16=79
4x=63
x=63/4 = 15.75 feet